If (H,*) is a subgroup of (G,*), and a * b is an element of H, must a is an element of H and b is an element of H? Explain your answer.
1
Expert's answer
2020-10-25T18:48:47-0400

Let "(G,*)" be a group containing at least 2 elements, and "e" be the identity of "(G,*)". Let "H=\{e\}". Recall that "e^{-1}=e" and "e*e=e." Then "(H,*)" is a subgroup of "(G,*)". If "a\in G" and "a\ne e", then "a^{-1}\ne e" as well. Put "b=a^{-1}". Then "a*b=a*a^{-1}=e\in H", but "a\notin H" and "b\notin H".